# Early Detection of Forest Calamities in Homogeneous Stands - Deep Learning Applied to Bark-Beetle Outbreaks

### v1

@@ 1 Introduction
Predicting disturbances is challenging due to the unpredictable nature of insect induced calamities [7]. An early and spatially precise detection of potential bark beetle infection center is needed to mitigate an epidemic bark beetle reproduction [8, 9]. Accordingly, several studies stressed the hypothesis of green attack detection based on remote sensing data [10, 11, 12]. However, there is wide consensus that green attack detection is either unreliable or infeasible on large spatial scales [13, 14, 9]. Instead, our focus is on identifying anomalous patterns in forest health that may serve as early warning indicators of potential disturbances. A limitation of early disturbance detection is often the lack of labeled training data. Satellite Image Time Series (SITS) based anomaly detection is a often unsupervised and do not rely on labeled training samples to capture seasonal variations [15, 16]. Often SITS approaches, such as the BFAST algorithm, learn temporal patterns and detect deviations from the normal state [17]. Other approaches utilize deep learning such as Long Short-Term Memory neural networks, that remember temporal patterns in time series [18]. Most approaches dependent on extensive historical information for every observed time series and often monitor univariate time series, creating hyperspectral relationships only by calculating vegetation indices (VIs) [15, 16, 18].
@@ 4.2 LSTM
$f_{t}=\sigma\left(w_{f}h_{t-1}+u_{f}x_{t}+b_{f}\right)$ (3)
$i_{t}=\sigma\left(w_{i}h_{t-1}+u_{i}x_{t}+b_{i}\right)$ (4)
$o_{t}=\sigma\left(w_{o}h_{t-1}+u_{o}x_{t}+b_{o}\right)$ (5)
$\tilde{C}_{t}=tanh\left(w_{C}h_{t-1}+u_{C}x_{t}+b_{C}\right)$ (6)
@@ 64 LSTM Units $\left(window\_size,64\right)$ 33,024
Consequently, we calculated the adjusted reconstruction errors, which are based on the MAE of different band combinations. This adjustment is intended to help to distinguish between healthy and unhealthy time points. We selected the following combinations:
@@ 5.3 Evaluation Framework
As mentioned in section 2, in remote sensing based forest health monitoring, labeled data is extremely difficult to get. It is not possible to determine the exact start of the bark beetle infestation from remote sensing [9]. Therefore, evaluating models that detect early anomalies to predict forest disturbance is very difficult if there is no labeled data. We classified defoliation in pixels based on VI thresholds and created a scoring method that assumes an anomaly window around the defoliation date. The earlier anomalies are detected within the window, the higher they get rewarded.
@@
where the difference between time of anomaly detection and defoliation time ($t_{\text{anomaly}}-t_{\text{defoliation}}$) is denoted as $lag$ and $k$ represents the anomaly window (52)
@@
If multiple anomalies were detected within the anomaly window, only the first anomaly was considered, while subsequent detections were assigned a score of zero to ensure that models are not rewarded multiple times for a single defoliation event.
The Pre-Defoliation Score provides a robust framework for evaluating early forest disturbance detection. By prioritizing timely anomaly identification while penalizing false alarms and delayed detections, this metric offers a meaningful way to compare models, even in the absence of perfectly labeled ground truth data. Future studies can leverage this approach to improve remote sensing-based monitoring of forest health.
@@ 5.4 Model Scoring
To assess the performance advantages of our models in comparison to established methods, we compared the LSTM-AE models to the break detection algorithm BFAST. BFAST detects structural breaks in time series trends, making it particularly suitable for analyzing VIs. We tested the BFAST algorithm on the following VIs: $NDVI$, $DWSI$, $EVI$, and $DRS$, which are frequently used in bark-beetle detection [15, 16, 61, 19, 62]. BFAST requires a historical reference period for each observed pixel, we set the training phase from January 2018 to January 2020 to ensure a sufficient number of observations. The algorithm outputs a detected break time for each time series, which we then use to compute the Pre-Defoliation Score, along with the performance metrics precision, accuracy and F1-score. Finally, we compared these results to those of the LSTM-AE models to evaluate their relative effectiveness in early anomaly detection.
@@ 6.3 Evaluation Maps
The Pre-Defoliation Scoring system allows enables the creation of maps that visualize model performance within an observed area. Fig. 14 visualizes the Pre-Defoliation Scores of the respective pixels in a larger spruce cultivation area for the time span from 2020 to the end of 2024.

### v2

@@ 1 Introduction
Predicting disturbances is challenging due to the unpredictable nature of insect induced calamities [7]. An early and spatially precise detection of potential bark beetle infection center is needed to mitigate an epidemic bark beetle reproduction [8, 9]. Accordingly, several studies stressed the hypothesis of green attack detection based on remote sensing data [10, 11, 12]. However, there is wide consensus that green attack detection is either unreliable or infeasible on large spatial scales [13, 14, 9]. Instead, our focus is on identifying anomalous patterns in forest health that may serve as early warning indicators of potential disturbances. A limitation of early disturbance detection is often the lack of labeled training data. Satellite Image Time Series (SITS) based anomaly detection is a often unsupervised and does not rely on labeled training samples to capture seasonal variations [15, 16]. Often SITS approaches, such as the BFAST algorithm, learn temporal patterns and detect deviations from the normal state [17]. Other approaches utilize deep learning such as Long Short-Term Memory neural networks, that remember temporal patterns in time series [18]. Most approaches dependent on extensive historical information for every observed time series and often monitor univariate time series, creating hyperspectral relationships only by calculating vegetation indices (VIs) [15, 16, 18].
@@ 4.2 LSTM
$f_{t}=\sigma\left(w_{f}x_{t}+u_{f}h_{t-1}+b_{f}\right)$ (3)
$i_{t}=\sigma\left(w_{i}x_{t}+u_{i}h_{t-1}+b_{i}\right)$ (4)
$o_{t}=\sigma\left(w_{o}x_{t}+u_{o}h_{t-1}+b_{o}\right)$ (5)
$\tilde{C}_{t}=tanh\left(w_{\tilde{C}}x_{t}+u_{\tilde{C}}h_{t-1}+b_{\tilde{C}}\right)$ (6)
@@ 64 LSTM Units $\left(window\_size,64\right)$ 33,024
Consequently, we calculated the adjusted reconstruction errors, which are based on the MAE of different band combinations. This adjustment is intended to help distinguishing between healthy and unhealthy time points. We selected the following combinations:
@@ 5.3 Evaluation Framework
As mentioned in section 2, in remote sensing-based forest health monitoring, labeled data is extremely difficult to obtain. The exact start of the bark beetle infestation cannot be determined from remote sensing [9]. Therefore, evaluating models that detect early anomalies to predict forest disturbance is very difficult if there is no labeled data. We classified defoliation in pixels based on VI thresholds and created a scoring method that assumes an anomaly window around the defoliation date. The earlier anomalies are detected within the window, the higher they get rewarded.
@@
where the difference between time of anomaly detection and defoliation time ($t_{\text{anomaly}}-t_{\text{defoliation}}$) is denoted as $lag$ and $k$ represents the anomaly window (52). A shift, denoted as $s$ is added to award an anomaly detected at the time of obvious defoliation with a positive value.
@@
If multiple anomalies were detected within the anomaly window, only the first anomaly was considered, while subsequent detections were assigned a score of zero to ensure that the models are not rewarded multiple times for a single defoliation event.
The Pre-Defoliation Score provides a robust framework to evaluate early forest disturbance detection. By prioritizing timely anomaly identification while penalizing false alarms and delayed detections, this metric offers a meaningful way to compare models, even in the absence of perfectly labeled ground-truth data. Future studies can leverage this approach to improve remote sensing-based forest health monitoring.
@@ 5.4 Model Scoring
To assess the performance advantages of our models compared to established methods, we compared the LSTM-AE models to the break detection algorithm BFAST. BFAST detects structural breaks in time series trends, making it particularly suitable for analyzing VIs. We tested the BFAST algorithm on the following VIs: $NDVI$, $DWSI$, $EVI$, and $DRS$, which are frequently used in bark-beetle detection [15, 16, 61, 19, 62]. BFAST requires a historical reference period for each observed pixel, we set the training phase from January 2018 to January 2020 to ensure a sufficient number of observations. The algorithm outputs a detected break time for each time series, which we then use to compute the Pre-Defoliation Score, along with the performance metrics precision, accuracy and F1-score. Finally, we compared these results to those of the LSTM-AE models to evaluate their relative effectiveness in early anomaly detection.
@@ 6.3 Evaluation Maps
The Pre-Defoliation Scoring system enables the creation of maps that visualize model performance within an observed area. Fig. 15 visualizes the Pre-Defoliation Scores of the respective pixels in a larger spruce cultivation area for the time span from 2020 to the end of 2024.